OpenStudy (anonymous):

What polynomial has roots of -5, -4, and 1 ? x3 - 8x2 - 11x + 20 x3 - x2 + 22x + 40 x3 + x2 - 22x - 40 x3 + 8x2 + 11x - 20

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

The polynomial that has roots m, n, p, is (x - k)(x - n)(x - p)

OpenStudy (mathstudent55):

Your roots are -5, -4, and 1. You write 3 binomials, each one being x minus the root. Then multiply them together to find the polynomial.

OpenStudy (mathstudent55):

Notice that your first 2 roots are negative and the binomials are written with x MINUS the root.

OpenStudy (anonymous):

So -5 times -4 times 1?

OpenStudy (mathstudent55):

First write out the 3 binomials. Then multiply them together.

OpenStudy (mathstudent55):

A binomial has to have x in it.

OpenStudy (mathstudent55):

x - (-5) x - (-4) x - 1 The first two binomials above can be simplified since each one has two negative signs.

OpenStudy (anonymous):

So then they change to x + 5, x + 4, and x - 1?

OpenStudy (mathstudent55):

Great. So far so good. Now you need to multiply them together. (x + 5)(x + 4)(x - 1)

OpenStudy (mathstudent55):

Before you multiply them together, you can save some time by looking at the constant term of each binomial. The product of these 3 binomials is going to be of the form: Ax^3 + Bx^2 + Cx + D The D is a constant term (number only, no variable) It comes from the product of 5, 4 and -1. What is 5 * 4 * (-1) = ?

OpenStudy (anonymous):

Ok so, x^3 +8x^2 -11 + 20?

OpenStudy (mathstudent55):

Once again, what is 5 * 4 * (-1) = ?

OpenStudy (anonymous):

-20*

OpenStudy (mathstudent55):

That's better. Since only one of the choices has -20 as the constant term, that has to be the answer.

OpenStudy (anonymous):

Perfect! Thank you! I hate to be a bother but could you help me with another? You make me understand more than my teacher

OpenStudy (mathstudent55):

We can check by multiplying, of course: (x + 5)(x + 4)(x - 1) = = (x^2 + 5x + 4x + 20)(x - 1) = (x^2 + 9x + 20)(x - 1) = x^3 + 9x^2 + 20x - x^2 - 9x - 20 = x^3 + 8x^2 + 11x - 20

OpenStudy (mathstudent55):

Sure enough, the constant term is -20.

OpenStudy (mathstudent55):

Go ahead, but pls start anew post. I'll help you there.

OpenStudy (anonymous):

Got it!